# Impact of Calibration and Position Errors on Astrophysical Parameters of the Hi 21cm Signal

### v1

@@ Abstract
The Epoch of Reionization (EoR) and Cosmic Dawn (CD) are pivotal stages during the first billion years of the universe, exerting a significant influence on the development of cosmic structure. The detection of the redshifted 21-cm signal from these epochs is challenging due to the dominance of significantly stronger astrophysical foregrounds and the presence systematics. This work used the 21cm E2E pipeline, followed by simulation methodology described [1] to conduct synthetic observations of a simulated sky model that includes both the redshifted 21-cm signal and foregrounds. A framework was constructed using Artificial Neural Networks (ANN) and Bayesian techniques to directly deduce astrophysical parameters from the measured power spectrum. This approach eliminates the need for explicit telescope effects correction in interferometric arrays such as SKA-Low and HERA. The present work investigates the impact of calibration and position errors on retrieving the redshifted 21-cm power spectrum for the above arrays. We assessed the effects of these inaccuracies on the deduced astrophysical parameters and established acceptable tolerance levels. Based on our results, the calibration error tolerance for ideal signal detection is 0.001 %. However, if the position errors exceed 0.048 arcseconds, the remaining foregrounds would obscure the target signal.
@@ 1 Introduction
The observation of Hi 21cm power spectrum using large interferometric arrays currently holds the most significant potential to observe the redshifted Hi 21cm line [12, 13, 14, 15]. The 21cm signal inherently encodes information about the underlying dark matter distribution and the properties of the ionizing sources. As a result, it has the potential to trace the history of reionization, reflecting the evolution of the average ionization state of the intergalactic medium (IGM) with redshift during the EoR. Detection of the redshifted Hi signal is the key science goal of several ongoing and future experiments. Recently, the detection of a global 21cm signal reported by the Experiment to Detect Global Epoch of Reionization Signature (EDGES) team [16]. However, the detection has been challenged by another independent experiment, SARAS [17, 18]. Besides EDGES and SARAS, there are other independent single antenna experiments like BIGHORNS [19], SCI-HI [20], and LEDA [21] are also aiming to detect the global signal but are yet to report any detection. Conversely, interferometers focusing on statistical fluctuations have yielded significant upper limits on the EoR power spectrum (PS) amplitude. The most sensitive operational interferometers, such as the GMRT, MWA, LOFAR, and HERA, have all established upper limits on the power spectrum amplitude of the signal [22, 23, 24, 25, 26, 27], but there has been no confirmed detection of the cosmological Hi 21cm signal.
@@
However, observing the 21cm signal is highly challenging due to bright astrophysical foregrounds [13, 28, 29, 30, 31, 32, 33, 34]. Additionally, other sources of contamination, such as the Earth’s ionosphere and instrument systematics, make detection even more difficult. The observation of the 21cm signal heavily relies on the accuracy of foreground removal and the use of instruments with high sensitivity and controlled systematics. Over the past decade, several novel methods have been proposed to quantify and mitigate each type of contamination for foreground avoidance and removal [35, 36, 37, 38, 39, 40, 41, 42] , model the systematics of the instruments [43, 44, 45, 46] , address calibration effects [47, 48, 49, 50, 51, 52, 53, 54], and account for the impact of the Earth’s ionosphere [55, 56, 57]. These advancements are paving the way for highly sensitive next-generation interferometers, such as the Hydrogen Epoch of Reionization Array (HERA, [58]) and the Square Kilometer Array (SKA-Low, [59]), to detect the 21-cm signal and characterize the multi-redshift power spectrum (PS). This will result in tighter constraints on astrophysical parameters in the early universe. The forthcoming SKA-Low is specifically designed to have the sensitivity needed to detect the PS precisely and is anticipated to generate tomographic images of the HII regions [60].
@@
The EoR signal can be distinguished from foreground contamination because the EoR signal exhibits a spectral structure and is inherently uniform in spatial wavenumber (k) space, whereas foregrounds are spectrally smooth [61, 62, 63, 64, 65, 66, 67]. The foregrounds’ smooth spectral nature, along with the inherent chromaticity of the instruments, limit the contamination to the ’wedge’ in cylindrical Fourier space (i.e., the 2D power spectrum). The area outside this wedge, where the foregrounds are less prominent than the EoR signal, is referred to as the ’EoR window.’ However, the interaction between astrophysical foregrounds and the instrument results in the leakage of wedge power into the clean modes of the window, a phenomenon known as ’mode mixing’. This mode mixing impacts the ”EoR window,” the region outside the wedge, complicating the detection process [68]. One major recurring challenge in detecting cosmic signals is improper calibration. In radio astronomy sky-based calibration is commonly used, but CD/EoR observations often produce inaccurate models due to low angular resolutions and noise confusion, resulting in residual errors and hindering target cosmological signal detection [69, 35, 48, 50]. The redundant calibration method, investigated by observatories such as HERA, addresses this by repeatedly measuring the redundant baselines of interferometers to correct for the incoming sky signal and instrumental parameters [70, 71].
@@
The signal-to-noise ratio (SNR) is a crucial factor in detecting faint cosmic signals. [69, 1] suggests that even small calibration errors of 0.1 % can significantly reduce the dynamic range and obscure the signal. Further investigation is needed to determine the tolerance of imperfections, which may lead to noise that obscures the detection of weak signals. This study aims to quantify the necessary tolerance for successfully detecting the redshifted 21-cm signal along with astrophysical parameter from cosmic reionization using various sensitive telescopes. To achieve this, we utilize a hybrid machine learning (ML) approach that integrates artificial neural networks (ANN) with Bayesian methods to analyze the effects of imperfections on observational data and their direct impact on the associated astrophysical parameters.
@@
In this work, we have developed an emulator using an artificial neural network (ANN) framework. This trained ANN emulator was employed as model statistics to constrain EoR astrophysical parameters from the total observed sky power spectrum, which includes the Hi signal and systematic effects, via a Bayesian inference process. The motivation for building ANN emulators arises from the computational challenges in EoR 21cm cosmology. Generating numerous observable model signals for multi-dimensional parameter space to perform Bayesian inference using semi-numerical or radiative transfer methods is computationally intensive. Additionally, incorporating telescope effects through simulated observations further increases the computational expense. To address these challenges, we adopted a formalism already utilized by several groups [75, 76], using emulators for EoR signal modeling instead of actual simulations. To construct the training datasets for the emulators, we performed simulated radio interferometric observations using a 21cm E2E pipeline with two different array configurations: SKA Low core and HERA. This allowed us to calculate the total observed sky power spectrum with telescope effect. We also studied the systematic biases introduced in the observed power spectrum by comparing it with the true power spectrum. Furthermore, we examined how these biases influence parameter extraction and quantified the tolerance levels for calibration and position errors necessary for the successful detection of the Hi 21cm signal from the EoR using these sensitive telescopes through Bayesian inferences. This is followup work by [1], which quantified the tolerance for calibration and position errors in detecting the Hi 21cm signal using various interferometers by analyzing variations in RMS in the image plane and visually examining the 2D and 1D power spectra. Our research investigates the effects of gain calibration and position errors on astrophysical parameters, quantifying the tolerance levels needed to obtain inferred parameters that closely match the true values. Specifically, this work highlights the impact of these errors on the 21cm signal’s astrophysical parameters when no mitigation techniques are applied to the residual foreground contamination caused by them. Furthermore, we are developing a mitigation pipeline, as outlined by Beohar E. et al., in prep, to effectively correct these errors, enabling the accurate inference of the true power spectrum and astrophysical parameters.
@@ 2.1 Hi 21cm Maps
We simulated lightcones in a comoving box with of grid cells for a range for redshift extent of . The constructed lightcone volume is saved on a grid of size 232 × 232 × 64 and projected on a World Coordinate System (WCS) to generate an input signal for simulating observations. For this study, we simulate different lightcones to probe different realizations of reionization history using varying EoR parameters. To simulate these lightcones, we use three key EoR parameters , , which can be tweaked to create different reionization histories. We describe these in detail below following [81]:
@@ 2.2 Foreground Models
The foreground component employed in this study specifically includes compact sources. Two distinct catalogues were utilized: the 400 MHz catalogue derived from the uGMRT observations of the ELAIS N1 field [33] and the Tiered Radio Extragalactic Continuum Simulation (T-RECS, [96]). The sources were repositioned around a declination of -30∘. Flux values were transformed to 142 MHz using , with an value of -0.8. The 400 MHz catalog ranged from 100 Jy to 0.6 Jy. T-RECS, simulating the continuum radio sky from 150MHz to 20GHz, models Active Galactic Nuclei (AGNs) and Star-Forming Galaxies (SFGs), incorporating real data for realistic cosmological evolution. The chosen catalogue for this study, covering approximately 3 deg2, was derived from the 25 deg2 T-RECS catalogue, maintaining a flux cut-off similar to the EN1 catalogue. Fluxes at 150 MHz were adjusted to 142 MHz using -0.8, resulting in a catalogue comprising 2522 sources within the selected field of view. Notably, the current analysis excludes a diffuse emission model in the foreground, aiming for simplicity, with future studies planning to integrate diffuse emission models to explore their impact.
@@ 3 Synthetic Observations
To conduct synthetic observations, we utilized a 21cm end-to-end (E2E) pipeline [1]. This pipeline employs the OSKAR software package [97] for simulating SKA Low configurations, while the Common Astronomy Software Application (CASA) [98] is used for simulating HERA configurations. In this simulated observation, the sky was observed with a phase center at and for a duration of 4 hours ( hours hour angle). The observing bandwidth of the lightcone spans 8 MHz with a channel separation of 125 kHz. For additional details, refer to Table 1.
@@ 3.1 Telescope Model
In this work, we are mainly focused on three different types of telescope model details are described below :
3.1.1 SKA Low
The Square Kilometre Array (SKA)222https://www.skao.int/en/explore/telescopes/ska-low is a next-generation, highly sensitive radio telescope designed to detect low-frequency radio signals, such as the 21-cm signal, making it a powerful tool for probing the Epoch of Reionization (EoR) [59]. It is expected to have sufficient sensitivity, with a Signal-to-Noise Ratio (SNR) greater than 1, to create tomographic images of the 21-cm field [60]. The SKA-Low array design includes 512 stations (referred to as AA4), each with a 35-meter diameter, comprising 256 antennas per station. Approximately 50 % of these stations will be concentrated within a 1 km diameter central core, while the remaining stations are distributed along three spiral arms, arranged in clusters of 6 stations with logarithmic spacing. The configuration allows for a maximum baseline of about 73.4 km. In this study, we focus on the SKA-Low layout, which consists of 307 stations with a maximum baseline of 73.4 km. From this layout, we selected all stations located within a 2 km radius (equivalent to a maximum baseline of 2000 meters from the central station), resulting in a subset of 231 stations, array shown in Fig 1.
3.1.2 HERA
The Hydrogen Epoch of Reionization Array (HERA)333https://reionization.org/ is an interferometer designed to detect fluctuations in the 21cm signal from the Cosmic Dawn and Epoch of Reionization [58]. Located in the South African Karoo Radio Astronomy Reserve, HERA will eventually consist of 350 parabolic dishes, each 14 meters in diameter, arranged in a highly redundant configuration—320 in a dense core and 30 as outriggers—as shown in Fig. 1. The instrument operates within the 100–200 MHz frequency range. This study employs the full array of 350 dishes, which provides a maximum baseline of approximately 880 meters.
@@ 3.1.2 HERA
Central Frequency | 142 MHz (z9) | 142 MHz (z9) | |
Bandwidth | 8 MHz | 8 MHz | |
Number of frequency channels | 64 | 64 | |
Field of view | 4∘ | 4∘ | |
Number of array elements () | 231 | 350 | |
Maximum baseline (m) | 2000 | 880 | |
Effective collective area () | 962 | 154 | |
Core area of an array () | 12.57 | 0.608 |
@@ 4 Power Spectrum
A key scientific objective of both ongoing and upcoming radio interferometric arrays is to detect and characterize fluctuations in the brightness temperature of the redshifted 21cm signal from the EoR. Various statistical tools are available for these interferometric experiments to calculate these fluctuations in terms of the power spectrum (PS) using both the visibility data [25] and the image plane [26]. In this study, we have calculated the theoretical PS from the image plane and the observed PS from simulated visibility data for different arrays such as SKA Low and HERA, as shown in Fig. 2.
@@ 4.2 Observational Power Spectrum
The interferometric visibility is defined as the correlation between the signals received by a pair of antennas, which is given by [100] :
Here, represents the baseline vector, while and correspond to the specific intensity and the antenna beam pattern, respectively, both as functions of frequency (). The term denotes the instrumental bandpass response. The unit vector is defined as , where , , and are the direction cosines pointing towards the east, north, and zenith, respectively, with . The solid angle element is given by . In this study, is assumed to be 1, meaning the primary beam effect is not taken into account.
@@
Here, represents the rest-frame frequency of the 21 cm spin-flip transition of Hi, while denotes the redshift corresponding to the observing frequency. The Hubble parameter is given by , and . The parameters and correspond to the matter and dark energy densities, respectively.

### v2

@@ Abstract
The Epoch of Reionization (EoR) and Cosmic Dawn (CD) are pivotal stages during the first billion years of the universe, exerting a significant influence on the development of cosmic structure. The detection of the redshifted 21-cm signal from these epochs is challenging due to the dominance of significantly stronger astrophysical foregrounds and the presence systematics. This work used the 21cm E2E (end to end) pipeline, followed by simulation methodology described [1] to conduct synthetic observations of a simulated sky model that includes both the redshifted 21-cm signal and foregrounds. A framework was constructed using Artificial Neural Networks (ANN) and Bayesian techniques to directly deduce astrophysical parameters from the measured power spectrum. This approach eliminates the need for explicit telescope layout effects correction in interferometric arrays such as SKA-Low. The present work investigates the impact of gain calibration and sky model position errors on the recovery of the redshifted 21-cm power spectrum for the SKA-Low AA∗ array configuration. We assessed the effects of these inaccuracies on the deduced astrophysical parameters and established acceptable tolerance levels. Based on our results, the gain calibration error tolerance for ideal signal detection is 0.001 %. However, if the sky model position errors exceed 0.048 arcseconds, the remaining foregrounds would obscure the target signal.
@@ 1 Introduction
The observation of Hi 21cm power spectrum using large interferometric arrays currently holds the most significant potential to observe the redshifted Hi 21cm line [12, 13, 14, 15]. The 21cm signal inherently encodes information about the underlying dark matter distribution and the properties of the ionizing sources. As a result, it has the potential to trace the history of reionization, reflecting the evolution of the average ionization state of the intergalactic medium (IGM) with redshift during the EoR. Detection of the redshifted Hi signal is the key science goal of several ongoing and future experiments. Recently, the detection of a global 21cm signal reported by the Experiment to Detect Global Epoch of Reionization Signature (EDGES) team [16]. However, the detection has been challenged by another independent experiment, SARAS [17, 18]. Besides EDGES and SARAS, there are other independent single antenna experiments like BIGHORNS [19], SCI-HI [20], and LEDA [21] are also aiming to detect the global signal but are yet to report any detection. Conversely, interferometers focusing on statistical fluctuations have yielded significant upper limits on the EoR power spectrum (PS) amplitude. The most sensitive operational interferometers, such as the GMRT, MWA, LOFAR, and HERA, have all established upper limits on the power spectrum amplitude of the signal [22, 23, 24, 25, 26], but there has been no confirmed detection of the cosmological Hi 21cm signal.
@@
However, observing the 21cm signal is highly challenging due to bright astrophysical foregrounds primarily Galactic synchrotron emission and extragalactic point sources that are several orders of magnitude brighter than the signal of interest [27, 28, 29, 30, 31, 32, 33]. Additionally, other sources of contamination, such as the Earth’s ionosphere and instrument systematics, make detection even more difficult. The observation of the 21cm signal heavily relies on the accuracy of foreground removal and the use of instruments with high sensitivity and controlled systematics. Over the past decade, several novel methods have been proposed to quantify and mitigate each type of contamination for foreground avoidance and removal [34, 35, 36, 37, 38, 39, 40, 41] , model the systematics of the instruments [42, 43, 44, 45] , address calibration effects [46, 47, 48, 49, 50, 51, 52, 53], and account for the impact of the Earth’s ionosphere [54, 55, 56]. These advancements are paving the way for highly sensitive next-generation interferometers, such as the Hydrogen Epoch of Reionization Array (HERA, [57]) and the Square Kilometer Array (SKA-Low, [58]), to detect the 21-cm signal and characterize the multi-redshift power spectrum (PS). This will result in tighter constraints on astrophysical parameters in the early universe. The forthcoming SKA-Low is specifically designed to have the sensitivity needed to detect the PS precisely and is anticipated to generate tomographic images of the HII regions [59].
@@
The EoR signal can be distinguished from foreground contamination because the EoR signal exhibits a spectral structure and is inherently uniform in spatial wavenumber (k) space, whereas foregrounds are spectrally smooth [60]. The foregrounds’ smooth spectral nature, along with the inherent chromaticity of the instruments, limit the contamination to the ’wedge’ in cylindrical Fourier space (i.e., the 2D power spectrum) [34, 35, 37]. The area outside this wedge, where the foregrounds are less prominent than the EoR signal, is referred to as the ’EoR window’ [61]. However, the interaction between astrophysical foregrounds and the instrument results in the leakage of wedge power into the clean modes of the window, a phenomenon known as ’mode mixing’ [61]. This mode mixing impacts the ”EoR window,” the region outside the wedge, complicating the detection process. One major recurring challenge in detecting cosmic signals is improper calibration. In radio astronomy sky-based calibration is commonly used, but CD/EoR observations often produce inaccurate models due to low angular resolutions and noise confusion, resulting in residual errors and hindering target cosmological signal detection [62, 34, 63, 49]. The redundant calibration method, investigated by observatories such as HERA, addresses this by repeatedly measuring the redundant baselines of interferometers to correct for the incoming sky signal and instrumental parameters [64]. However, as demonstrated by [65], redundant calibration remains susceptible to errors introduced during the absolute calibration step, which necessarily depends on a sky model. Even in the case of a perfectly redundant array with identical antenna beams, incompleteness in the sky model leads to frequency-dependent calibration errors that can contaminate the 21cm power spectrum. To mitigate these limitations, [66] introduced a unified Bayesian calibration framework that integrates both sky-based and redundant approaches. This framework explicitly accounts for instrumental systematics, such as antenna position offsets, beam non-uniformities, and incomplete sky models, thereby improving the robustness and accuracy of 21cm power spectrum estimation.
@@
The signal-to-noise ratio (SNR) is a critical factor in the detection of faint cosmic signals, such as the redshifted 21-cm line from the epoch of reionization. [34] demonstrated that even minor calibration errors can significantly reduce the dynamic range, thereby obscuring the cosmological signal. Building on this, [1], through an analysis of the one-dimensional power spectrum (1D PS), determined that the optimal calibration error tolerance for reliable signal detection is approximately . Consistently, [63] showed that calibration errors must be limited to below approximately or 0.001 % in amplitude in order to prevent contamination of the EoR window in power spectrum measurements. Despite these insights, further investigation is needed to determine the precise tolerance limits for various instrumental imperfections that could hinder weak signal detection. This study aims to quantify the tolerance levels required for the successful detection of the redshifted 21-cm signal and the recovery of astrophysical parameters from the epoch of reionization using the SKA-Low telescope. To achieve this, we utilize a hybrid machine learning (ML) approach that integrates artificial neural networks (ANN) with Bayesian methods to analyze the effects of imperfections on observational data and their direct impact on the associated astrophysical parameters.
@@
In this work, we have developed an emulator using an artificial neural network (ANN) framework. This trained ANN emulator was employed as model statistics to constrain EoR astrophysical parameters from the total observed sky power spectrum, which includes the Hi signal and systematic effects, via a Bayesian inference process. The motivation for building ANN emulators arises from the computational challenges in EoR 21cm cosmology. Generating numerous observable model signals for multi-dimensional parameter space to perform Bayesian inference using semi-numerical or radiative transfer methods is computationally intensive. Additionally, incorporating telescope layout effects through simulated observations further increases the computational expense. To address these challenges, we adopted a formalism already utilized by several groups [70, 71], using emulators for EoR signal modeling instead of actual simulations. To construct the training datasets for the emulators, we performed simulated radio interferometric observations using a 21cm E2E pipeline for SKA-Low core array configurations. This allowed us to calculate the total observed sky power spectrum with telescope layout effect. We also studied the systematic biases introduced in the observed power spectrum by comparing it with the true power spectrum. Furthermore, we examined how these biases influence parameter extraction and quantified the tolerance levels for calibration and position errors necessary for the successful detection of the Hi 21cm signal from the EoR using this sensitive telescope through Bayesian inferences. This is followup work by [1], which quantified the tolerance for calibration and position errors in detecting the Hi 21cm signal using various interferometers by analyzing variations in RMS in the image plane and visually examining the 2D and 1D power spectra. Our research investigates the effects of gain calibration and position errors on astrophysical parameters, quantifying the tolerance levels needed to obtain inferred parameters that closely match the true values. Specifically, this work highlights the impact of these errors on the 21cm signal’s astrophysical parameters when no mitigation techniques are applied to the residual foreground contamination caused by them. Furthermore, we are developing a mitigation pipeline, as outlined by Beohar E. et al., in prep, to effectively correct these errors, enabling the accurate inference of the true power spectrum and astrophysical parameters.
@@ 2.1 Hi 21cm Maps
We simulated lightcones in a comoving box with of grid cells for a range for redshift extent of . The constructed lightcone volume is saved on a grid of size 232 × 232 × 64 and projected on a World Coordinate System (WCS) to generate an input signal for simulating observations. For this study, we simulate different lightcones to probe different realizations of reionization history using varying EoR parameters. To simulate these lightcones, we use three key EoR parameters , , and which can be tweaked to create different reionization histories. These parameters are commonly used to characterize the EoR, as they effectively capture the timing, source population, and morphology of reionization while remaining physically interpretable and computationally efficient [85]. We describe these in detail below following [85, 76]:
@@ 2.2 Foreground Models
The foreground component employed in this study includes compact sources based on the Tiered Radio Extragalactic Continuum Simulation (T-RECS, [92]). T-RECS simulates the continuum radio sky from 150 MHz to 20 GHz, modeling Active Galactic Nuclei (AGNs) and Star-Forming Galaxies (SFGs), incorporating observational constraints for realistic cosmological evolution. For this study, a subset covering approximately 4 deg2 was extracted from the full 25 deg2 T-RECS catalogue, with a flux cut-off applied between 0.6 Jy to 3.1 mJy at 150 MHz. Flux densities originally specified at 150 MHz were extrapolated to 142 MHz using a spectral index , yielding a total of 2522 compact sources within the chosen field of view. To maintain simplicity, the current analysis focuses solely on compact sources and excludes both diffuse foreground emission and complex sources with extended morphology or non-power-law spectral characteristics. Future studies will incorporate these additional components to evaluate their influence on signal recovery.
@@ 3 Synthetic Observations
To conduct synthetic observations, we utilized a 21cm end-to-end (E2E) pipeline [1]. This pipeline employs the OSKAR software package [93] for simulating SKA-Low configurations, the Common Astronomy Software Applications (CASA) package [94] for further reading and processing of the visibility data. In this simulated observation, the sky was observed with a phase center at and for a duration of 4 hours ( hours hour angle). The observing bandwidth of the lightcone spans 8 MHz with a channel separation of 125 kHz. For additional details, refer to Tab. 1.
@@ 3.1 Telescope Model
In this work, we focus primarily on the upcoming SKA-Low telescope. The Square Kilometre Array (SKA)222https://www.skao.int/en/explore/telescopes/ska-low is a next-generation, highly sensitive radio telescope designed to detect low-frequency radio signals, such as the 21-cm signal, making it a powerful tool for probing the Epoch of Reionization (EoR) [58]. The SKA-Low array design includes 512 stations (referred to as AA4), each with a 35-meter diameter, comprising 256 antennas per station. Approximately 50 % of these stations will be concentrated within a 1 km diameter central core, while the remaining stations are distributed along three spiral arms, arranged in clusters of 6 stations with logarithmic spacing. The configuration allows for a maximum baseline of about 73.4 km. In this study, we focus on the SKA-Low layout, which consists of 307 stations with a maximum baseline of 73.4 km. From this layout, we selected all stations located within a 2 km radius (equivalent to a maximum baseline of 2000 meters from the central station), resulting in a subset of 231 stations, array shown in Fig. 1.
@@
Central Frequency | 142 MHz (z9) |
Bandwidth | 8 MHz |
Number of frequency channels | 64 |
Image Field of view | 4∘ |
Number of array elements () | 231 |
Maximum baseline (m) | 2000 |
Effective collective area () | 962 |
Core area of an array () | 12.57 |
@@ 4 Power Spectrum
A key scientific objective of both ongoing and upcoming radio interferometric arrays is to detect and characterize fluctuations in the brightness temperature of the redshifted 21cm signal from the EoR. Various statistical tools are available in interferometric experiments to estimate the 21-cm brightness temperature fluctuations through the power spectrum (PS), either directly from the measured visibilities using the delay spectrum technique [95, 96], or via image-based methods that involve gridding, calibration, and deconvolution to reconstruct the full 2D or 3D power spectrum [97, 50]. In this study, we compute the theoretical power spectrum (PS) from the image plane and the observed PS from simulated visibility data using the SKA-Low layout, as illustrated in Fig. 2.
@@ 4.2 Observational Power Spectrum
The interferometric visibility is defined as the correlation between the signals received by a pair of antennas, which is given by [38, 99] :
Here, represents the baseline vector, while and correspond to the specific intensity and the antenna beam pattern, respectively, both as functions of frequency (). The term represents the instrumental bandpass response. For this simulated observation, we assume an isotropic, frequency-independent antenna beam pattern, resulting in a flat bandpass response with = 1. The unit vector is defined as , where , , and are the direction cosines pointing towards the east, north, and zenith, respectively, with . The solid angle element is given by . In this study, is assumed to be 1, meaning the primary beam effect is not taken into account.
@@
Here, represents the rest-frame frequency of the 21 cm spin-flip transition of Hi, while denotes the redshift corresponding to the observing frequency. The Hubble parameter is given by , and . The parameters and correspond to the matter and dark energy densities, respectively [100].